jigglypuff314 (jigglypuff314):

dy/dx = 4y/(x+3) I need to find the anti-derivative of that, pwease help :3

OpenStudy (anonymous):

divide both side by y and multiply both sides by dx

OpenStudy (anonymous):

then integrate both sides. Done

jigglypuff314 (jigglypuff314):

dy/y = 4dx / (x+3) ? so that becomes lny on the left but idk about the right side...

OpenStudy (anonymous):

it's just integrating 4/(x+3) dx

jigglypuff314 (jigglypuff314):

so lny = 4ln(x+3) ?

OpenStudy (anonymous):

lny = 4ln(x+3) + C

jigglypuff314 (jigglypuff314):

oh right plus C :P and how do I solve for y? ... divide both sides by ln?

OpenStudy (anonymous):

XDDDD!!! sorry i just have to laugh a little. you can't dividing both sides by just ln!! it's like dividing both sides by just a square root sign!!

jigglypuff314 (jigglypuff314):

:P yeah... I kinda might have missed the log lesson xD

OpenStudy (anonymous):

ln(y) = x means y = e^x

jigglypuff314 (jigglypuff314):

so y = e^(4ln(x+3) + C) ?

OpenStudy (anonymous):

right

jigglypuff314 (jigglypuff314):

y = e^(4ln(x+3)) * e^(c) = C e^(4ln(x+3)) ? what else?

OpenStudy (anonymous):

n ln(x) = ln(x^n)

jigglypuff314 (jigglypuff314):

so I should bring the 4 up like y = C e^(ln((x+3)^4)) ?

OpenStudy (anonymous):

and e^ln(x) = x

jigglypuff314 (jigglypuff314):

so y = C (x+3)^4 ?

OpenStudy (anonymous):

bingo

jigglypuff314 (jigglypuff314):

Alright :D Thank you!!! ^_^

OpenStudy (anonymous):

yw